Our sun has a wavelength of maximum intensity of emission approximately 510 nm.


b) Using the Stefan–Boltzmann law, and given that the luminosity of the sun is 3.9 x 10^26 W, calculate the radius of the sun, giving your answer to two significant figures.


1
Expert's answer
2021-04-19T07:22:47-0400

From the Wien's displacement law:


"\lambda_{max} =\dfrac{b}{T}"

where "\lambda_{max} = 510nm = 5.1\times 10^{-7}m" is the  wavelength of maximum intensity, "T" is the absolute temperature of the Sun, and "b = 2.9\times 10^{-3}m\cdot K" is the Wien's displacement constant. Thus, the temperature of the Sun is:


"T = \dfrac{b}{\lambda_{max}}"

According to the Stefan–Boltzmann law:


"\dfrac{L}{A} = \sigma T^4"

where "L = 3.9 \times 10^{26} W" is the luminosity of the Sun, "A" is the area of Sun's surface, "\sigma = 5.67\times 10^{-8}W/(m^2\cdot K^4)" is the Stefan–Boltzmann constant. Substituting the expression for "T" and expressing "A", obtain:


"A = \dfrac{L}{\sigma T^4} = \dfrac{L\lambda_{max}^4}{\sigma b^4}"

On the other hand, the area of a sphere is:


"A = \pi R^2"

where "R" is the radius of the Sun. Expressing R, find:


"R = \sqrt{\dfrac{A}{\pi}} = \sqrt{ \dfrac{L\lambda_{max}^4}{\pi\sigma b^4}}= \dfrac{\lambda_{max}^2}{b^2}\sqrt{\dfrac{L}{\pi\sigma}}\\
R = \dfrac{(5.1\times 10^{-7})^2}{(2.9\times 10^{-3})^2}\sqrt{\dfrac{3.9 \times 10^{26}}{\pi\cdot 5.67\times 10^{-8}}} \approx 1.45\times 10^{9}m"

Answer. "1.45\times 10^{9}m"